# Multilayer Network of Cardiovascular Diseases and Depression via Multipartite Projection

### v1

@@ Discussion
In this study, we proposed a multipartite projection method to infer the relationships between CVD and depression, leading to the creation of a weighted multilayer network of CVD-related phenotypes and depressive symptoms. The proposed method can be considered a generalization of previous disease network modeling approaches [13, 15, 16]. It demonstrates several advantages: (i) it uses MI correlation to capture both linear and nonlinear associations; (ii) it is capable of integrating multiple omics datasets into the projection; (iii) it defines a refined projection strategy that accounts for shared biomarkers considering their correlations in a weighted way. Additionally, this method has the potential to incorporate directed causal relationships and higher-order interactions (e.g. synergistic relationship) into the projection. It also has the capability to construct weighted higher-order projected disease networks by considering the shared neighboring nodes of more than two disease variables.
@@
We applied the method to the YFS dataset and inferred the projected scores between CVD-related phenotypes and depressive symptoms considering two available omics datasets: metabolome and lipidome. A weighted multilayer disease network was created based on the projected scores. Each layer of the projected network represents the links between diseases in a specific omics context, which is metabolomic or lipidomic in this study. This framework allows for a rich network representation of CVD and depression across multiple dimensions and facilitates the identification of related biomarkers. It provides better insights into possible causal mechanisms and biological pathways underlying the links between these diseases, getting closer to explaining comorbidities.
@@
The most contributing metabolites and lipids (sorted in Figure 4) are inferred to play a role in the biological pathways connecting CVD and depression. In metabolites, Creatinine is a byproduct of breaking down creatine. On the depression side, creatine presents an antidepressant-like effect [24, 25]. On the CVD side, creatinine is associated with several traditional cardiovascular risk factors such as diabetes and dyslipidemia in hypertensive patients and stroke [26]. Valine has been considered to be significantly associated with depression in recent studies [27, 28]. High levels of branched chain amino acids (BCAAs), which both valine and leucine are categorized as, predicted a higher risk of developing CVD [27, 29]. Such particles are also rich in phospholipids which are known to play a crucial role in protecting against cardiovascular disease by promoting cholesterol efflux and reducing inflammation [30, 31]. The anti-inflammatory effects of phospholipids within HDL may simultaneously improve depressed mood and increase treatment response to conventional antidepressant medications in patients with major depression [32]. Such an anti-inflammatory property is also associated with apolipoprotein A-I, a major protein component of HDL particles [33]. This may well explain why "ApoBApoA1" is also in the ranking. Further, high cholesterol has been recognized as one of the significant risk factors for cardiovascular disease (CVD) [34]. It also shows a positive correlation with depression severity [35].
@@
On the lipid side, the projection method identified lipids such as several specific SM, PC, TAG, DAG, PG and PE as significant biomarkers (Figure 4). PC and PE are two major classes of phospholipids that have been associated with CVD and depression. They play important roles in numerous biological processes and are the most abundant phospholipids in mammalian cell membranes. Their levels have been linked to CVD risk and outcomes [36]. Alterations in PC and PE, have been observed in individuals with psychotic experiences, which are associated with an increased risk of mental disorders, including depression [37]. Additionally, TAG and DAG have also been reported to be associated with both CVD and depression in many studies [38, 39, 40, 41].
@@
When turning our attention to risk factors, the projected network identified sex and BMI as two important risk factors linking CVD and depression. Sex is indeed a risk factor (e.g., covariate, confounder) for these health conditions [48, 12]. Men, on average, are more prone to certain risks, such as high blood pressure, higher levels of LDL (bad) cholesterol, and a higher likelihood of smoking and excessive alcohol consumption. Consequently, the risk of developing CVD tend to begin earlier in men, often around middle age. Conversely, women are more likely to be diagnosed with depression [49]. The sex-specific difference in depression can be attributed to the sex-induced biological changes such as in hormones in women, and different presentations of depressive symptoms [50]. It is important to note that sex can act as a covariate or confounder. Thus, we conducted gender-specific analyses using the projection method. Further discussion on these results is provided in S2.3 in the Supplementary Information.
@@
The biomarkers underlying the strong links between risk factors and phenotypes were detected. These biomarkers are considered important in elucidating how risk factors contribute to disease development and understanding the underlying biological pathways involved. For instance, creatinine is a gender-specific metabolite [56, 57], and has been recognized as a nutritional marker associated with decreased appetite [58, 59]. SM 32.2 and many specific LPCs are identified as biomarkers related to appetite in individuals with depression, although no human studies have reported this finding. However, a recent paper using mouse models discovered that an increase in LPCs in the blood following a prolonged fasting increased lysophosphatidic acid (LPA) level in the brain [60]. These LPAs can modulate the excitability of nerve cells in the cerebral cortex, thereby influencing food intake and appetite regulation. Additionally, the detection of TAG and DAG in the context of sex-related cardiovascular research suggests potential lipid-related mechanisms underlying sex disparities in CVD susceptibility and outcomes [61]. TAG concentrations are consistently lower in healthy lean and obese women than in men [62], and its elevated level has been reported to be highly associated with CVD [63, 64, 38].
@@
This study has several limitations. Firstly, we conducted two steps to eliminate redundant information based on MI correlation. However, this process might lead to an underestimation of projected scores, as some intermediate biomarkers that are not entirely redundant may have been omitted. Secondly, the YFS dataset includes only limited omics data: metabolome and lipidome, which may not be sufficient for a complete construction of the projected links. These two limitations likely contribute to the consistently low explaining score (see S2.1 in the Supplementary Information) for all projection definitions. This finding underscores the importance of including multiple omics data in the projection. Another limitation pertains to the absence of directionality in the MI correlation. MI is a correlation measure rather than a causal discovery method, which makes it incapable of capturing causal relationships between variables. It is therefore challenging to determine whether a biomarker or risk factor plays a role of a confounder, collider, mediator or moderator. Finally, our method is blind to so-called higher-order correlations or synergistic associations beyond pairwise relationships. This is a common limitation in the field, as similar approaches in the literature have the same constraint.
@@
This study may provide direction for future data gathering efforts towards biomarkers that are predicted as highly important unexpectedly such as specific SM and PG. Future work should also involve incorporating directed causal [65] and higher-order relationships [66] into the projection and replicating the analysis across diverse datasets, including more omics data such as genomics and proteomics.
@@ Conclusions
The findings of this study underscore the importance of a multidimensional approach in studying complex diseases. The multipartite projection method offers a powerful tool for uncovering the intricate biological mechanisms underlying comorbid conditions, paving the way for more targeted preventive and therapeutic strategies. Future research could expand this methodology to other disease pairs and incorporate additional omics datasets to further enhance our understanding of disease comorbidities.

### v2

@@ Discussion
In this study, we proposed a multipartite projection method to infer complex disease interactions between CVD and depression, resulting in the creation of weighted multilayer networks encompassing CVD-related phenotypes and depressive symptoms. The proposed method can be considered a generalization of previous disease network modeling approaches [13, 15, 16]. It demonstrates several key advantages by employing mutual information (MI) correlation and incorporating multi-omics datasets. First, MI captures both linear and nonlinear associations, providing a more comprehensive understanding of the interdependencies between variables. Second, incorporating multiple omics datasets into a single projection allows for a richer and more detailed analysis of the biological pathways involved. Third, the method introduces a refined projection strategy that accounts for shared biomarkers in a weighted manner, thereby offering a more nuanced approach to quantifying disease associations. Moreover, the proposed method extends beyond existing approaches by offering the potential to incorporate directed causal relationships and higher-order interactions, such as synergistic associations, into the projection. This capability is particularly important for understanding the complex and often non-linear interactions that may underlie comorbid conditions like CVD and depression. It also has the capability to construct multilayer higher-order disease networks by considering the shared neighboring nodes of more than two disease variables.
@@
We applied the method to the YFS dataset [Raitakari2008] and inferred the projected scores between CVD-related phenotypes and depressive symptoms considering two available omics datasets: metabolome and lipidome. A weighted multilayer disease network was created based on the projected scores. Each layer of the projected network represents the links between diseases in a specific omics context, which is metabolomic or lipidomic in this study. This framework allows for a rich network representation of CVD and depression across multiple dimensions and facilitates the identification of related biomarkers. It provides better insights into possible biological pathways and mechanisms underlying the projected links between these diseases, getting closer to explaining comorbidities.
@@
The metabolites and lipids identified as the most significant contributors (sorted in Figure 4) are inferred to play key roles in the biological pathways connecting CVD and depression. Among metabolites, creatinine (a byproduct of creatine metabolism) emerged as a significant factor. In the context of depression, creatine has antidepressant-like effects [24, 25]. Conversely, in the realm of CVD, elevated creatinine levels are associated with traditional cardiovascular risk factors such as diabetes, dyslipidemia in hypertensive patients, and stroke [26]. These findings can also be observed in our study, as detailed in Section S2.2 of the Supplementary Information. Similarly, valine, a branched-chain amino acid (BCAA), has been highlighted in recent studies for its significant association with depression [27, 28]. Elevated levels of BCAAs, including both valine and leucine, have been linked to an increased risk of developing CVD [27, 29]. These particles are also rich in phospholipids, which are known to be instrumental in protecting against cardiovascular disease by promoting cholesterol efflux and reducing inflammation [30, 31].
Phospholipids within high-density lipoprotein (HDL) particles are particularly noteworthy for their anti-inflammatory effects, which may not only protect against CVD but also improve mood and enhance the efficacy of conventional antidepressant treatments in patients with major depression [32]. This anti-inflammatory property is also associated with apolipoprotein A-I, a major protein component of HDL particles [33], which may explain the high ranking of variable "ApoBApoA1" in our analysis. Further, high cholesterol has long been recognized as one of the significant risk factors for CVD [34]. It also shows a positive correlation with depression severity [35].
@@
The projection method also identified several specific lipids, including sphingomyelins (SM), phosphatidylcholines (PC), triacylglycerols (TAG), diacylglycerols (DAG), phosphatidylglycerols (PG), and phosphatidylethanolamines (PE), as significant biomarkers (Figure 4). Among these, PC and PE are two major classes of phospholipids that are associated with CVD and depression. They play crucial roles in various biological processes, including those related to CVD risk and outcomes [36], and are the most abundant phospholipids in mammalian cell membranes. Alterations in PC and PE, have been observed in individuals with psychotic experiences, which are associated with an increased risk of mental disorders, including depression [37]. Additionally, TAG and DAG have also been reported to be associated with both CVD and depression in many studies [38, 39, 40, 41].
@@
When turning our attention to risk factors, the projected network identified sex and BMI as two important risk factors linking CVD and depression. Sex is indeed a risk factor (e.g., covariate, confounder) for these health conditions [48, 12]. Men, on average, are more prone to certain risks, such as high blood pressure, higher levels of LDL (bad) cholesterol, and a higher likelihood of smoking and excessive alcohol consumption. Consequently, the risk of developing CVD tend to begin earlier in men, often around middle age. Conversely, women are more likely to be diagnosed with depression [49]. The sex-specific difference in depression can be attributed to the sex-induced biological changes such as in hormones in women, and different presentations of depressive symptoms [50]. It is important to note that sex can act as a covariate or confounder. Thus, we conducted gender-specific analyses using the projection method. Further discussion on the results is provided in Section S2.3 of the Supplementary Information.
@@
The biomarkers underlying the strong links between risk factors and phenotypes were also identified. These biomarkers are considered important in elucidating how risk factors contribute to disease development and understanding the underlying biological pathways involved. For instance, creatinine is a gender-specific metabolite [56, 57], and has been recognized as a nutritional marker associated with decreased appetite [58, 59]. SM 32.2 and many specific LPCs are identified as biomarkers related to appetite in individuals with depression. Although no human studies have reported this finding, a recent paper using mouse models discovered that an increase in LPCs in the blood following a prolonged fasting increased lysophosphatidic acid (LPA) level in the brain [60]. These LPAs can modulate the excitability of nerve cells in the cerebral cortex, thereby influencing food intake and appetite regulation. Additionally, the detection of TAG and DAG in the context of sex-related cardiovascular research suggests potential lipid-related mechanisms underlying sex disparities in CVD susceptibility and outcomes [61]. TAG concentrations are consistently lower in healthy lean and obese women than in men [62], and its elevated level has been reported to be highly associated with CVD [63, 64, 38].
@@
This study has several limitations. Firstly, we conducted two steps to eliminate redundant information based on MI correlation. However, this process might lead to an underestimation of projected scores, as some intermediate biomarkers that are not entirely redundant may have been omitted. Secondly, the YFS dataset includes only limited omics data: metabolome and lipidome, which may not be sufficient for a complete construction of the projected links. These two limitations likely contribute to the consistently low explaining score (see Section S2.1 of the Supplementary Information) for all projection definitions. This finding underscores the importance of including multiple omics data in the projection. Another limitation pertains to the absence of directionality in the MI correlation. MI is a correlation measure rather than a causal discovery method, which makes it incapable of capturing causal relationships between variables. It is therefore challenging to determine whether a biomarker or risk factor plays a role of a confounder, collider, mediator or moderator. Finally, our method is blind to so-called higher-order correlations or synergistic associations beyond pairwise relationships. This is a common limitation in the field, as similar approaches in the literature have the same constraint.
@@
This study may provide direction for future data collection efforts towards biomarkers that are unexpectedly predicted as highly important, such as specific SM and PG. Future work should also involve incorporating directed causal [65] interactions and higher-order synergistic associations [66] into the projection and replicating the analysis across diverse datasets, including more omics data such as genomics and proteomics.
@@ Conclusions
The findings of this study underscore the importance of a multidimensional approach in studying complex disease interactions. The multipartite projection method offers a powerful tool for uncovering the intricate biological pathways and mechanisms underlying comorbid conditions, paving the way for more targeted preventive and therapeutic strategies. Future research could expand this methodology to other disease pairs and incorporate additional omics datasets to further enhance our understanding of disease comorbidities.